OpenStudy (anonymous):

Math help please?:/ so im doing a chapter review and theres a couple that i just cant figure out. multiply or divide 12. a^2 + 5a + 4/ a3 ÷ a^2 + 3a + 2 / a^2 - 2a 13. x^2 + 13x + 40/ x - 7 ÷ x + 8/ x^2 - 49 thanks so much to anyone thatll try to help^-^

OpenStudy (anonymous):

The thing is, whenever you divide two fractions, you take the reciprocal of the divisor fraction to change the operation sign to multiplication. Example - \[\frac{ 4 }{ 9 } \div \frac{ 7 }{ 6 } => \frac{ 4 }{ 9 } \times \frac{ 6 }{ 7 },\]

OpenStudy (anonymous):

thank you for trying to help^-^ number 13. was x + 5 x+ 7 and then number 12 was (a+4)(a - 2) / a^2 ( a +2 ) thank you so much though, you put me on the right track!

OpenStudy (mathstudent55):